# DynamicScaler: Seamless and Scalable Video Generation for Panoramic Scenes

### v1

@@ 3.2 Global Motion Guidance
Figure 6: Example frames from a panorama video at the 0th, 16th, 32th, 48th, 64th and 80th frames. Despite the increasing video length, the visual quality of the panorama remains consistent, demonstrating the effectiveness of our method in generating long videos.
3.4 Infinite Long Scene Video Generation
@@
A major challenge in dynamic scene synthesis is generating long-duration video sequences, as current video diffusion models typically produce only short clips. To address this, we extended Offset Shifting Denoising from spatial to temporal dimension and introduced a temporal shifting strategy that enables continuous motion and the generation of videos with much longer duration by regular video diffusion models that have 16 frames capacity. By temporally shifting the denoising process, we ensure smooth transitions between frames, as shown in Figure 6.
@@
Moreover, by applying similar looping mechanism in our spatially shifting denoising process to the temporal dimension, we can achieve seamless looping video generation. Specifically, the first and last frame are connected and windows are allowed to cross the start and end boundary, treating the video frames as a ring, to create a loopable sequence with smooth, continuous motion. For more details please refer to our appendix.
4 Experiments
4.1 Qualitative Results
@@
Source Tuning-Free Arbitrary Size 360° Field-of-View Text Only Condition Image Condition Video Length Loopable
360DVD [23] CVPR24 $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 16 $\times$
4K4DGen [17] Arxiv24 $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ 16 $\times$
Scalecrafter [10] ICLR24 $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ 16 $\times$
VividDream [14] Arxiv24 $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ 16 $\times$
DynamicScaler - $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Infinite $\checkmark$
@@
Image Quality $\uparrow$ Dynamic Degree $\uparrow$ Motion Smoothness $\uparrow$ Temporal Flickering $\uparrow$ Q-Align(V) $\uparrow$
@@
We evaluate the performance of DynamicScaler against state-of-the-art models. We detail the implementation, experimental settings, comparison methods, and user studies conducted to assess the efficacy of our approach. The proposed DynamicScaler is built on top of existing text-to-video generation model VideoCrafter2 [3] and image-to-video generation model DynamiCrafter [26].
Compared to previous methods, DynamicScaler supports a wider range of settings and can generate videos of infinite length, while other methods are limited to generating only finite-length videos. Our approach generates dynamic panoramas at various resolutions, including both standard and 360° Field-of-View (FoV) panoramas, with support for both text- and image-conditioned inputs, as shown in Table 1. Currently, among the available concurrent works, only 360DVD is capable of generating dynamic panoramas, so we compare our method directly with it. We assess the visual quality by randomly selecting views from 100 generated video cases using random camera positions. Key factors such as image quality, dynamic range, motion smoothness, temporal flickering, and scene richness are evaluated according to the protocols of VBench [11] and CogVideoX [27]. Furthermore, we incorporate an LLM-based visual evaluator, Q-Align [24], to score both image and video quality. As shown in Table 1, DynamicScaler outperforms 360DVD in all metrics, demonstrating superior video quality in handling complex dynamic scenes.
@@ 4.2 Quantitative Results
As demonstrated in Figure 4, our approach generates panoramas of arbitrary sizes, supporting rectangular and 360° Field of View (FoV) configurations, offering unparalleled flexibility for diverse applications. As shown in Figure 7, the left and right edges of each generated frame align seamlessly, ensuring flawless integration into a 360° view—an essential feature for immersive environments. This highlights the robustness of our shift denoise technique in both non-360° and 360° settings. Furthermore, our method facilitates the creation of scene-level, long duration, and loopable videos through temporal shifting, without requiring additional training or compromising visual quality.
@@ 4.3 User Studies
We conduct user studies to evaluate videos based on five criteria: graphics quality, frame consistency, left-right continuity, content distribution, and motion patterns. 20 Participants are asked to select the video with the highest quality. Table 2 shows that DynamicScaler performs favorably in all criteria. We use two comparison types: same-case comparison (using the same scene for direct comparisons) and random-case comparison (using different scenes due to limited publicly available data). DynamicScaler consistently receives the highest ratings, especially for continuity, content distribution, and motion patterns, demonstrating its superior ability to generate seamless, dynamic panorama.
Figure 7: This figure demonstrates the seamlessness of generated panorama videos by horizontally concatenating frames, showcasing continuity across the left and right boundaries.
@@ 4.4 Ablation Studies
We conducted ablation studies to evaluate the impact of each core component in DynamicScaler under four different configurations: Direct Inference, where the video is generated directly at the target resolution without enhancement techniques, but results in an out-of-memory (OOM) error due to high resource demands; Without OSD (Offset Spatial Denoising), which excludes the offset spatial denoising mechanism that mitigates multi-scale artifacts; Without GMG (Global Motion Guidance), where global motion guidance is omitted, reducing frame-to-frame motion continuity; and the Full Method, which integrates all components. As shown in Table 3, the Full Method consistently outperforms all other configurations across all evaluation metrics, underscoring the significance of each component. Both OSD and GMG play critical roles in improving image quality, dynamic range, motion smoothness, and temporal coherence, while also minimizing temporal flickering.
5 Conclusion
This paper presents DynamicScaler for scalable, coherent panoramic dynamic scene synthesis. By introducing a Spatial Rotating Denoiser and Seamless Rotating Window, our approach ensures efficient denoising and consistent boundary transitions. The Global Motion Guidance mechanism maintains local detail and global motion continuity, delivering superior content quality and motion smoothness. DynamicScaler outperforms existing methods in scalability and performance, offering a practical, training-free solution for creating high-quality, immersive AR/VR dynamic content across various resolutions and aspect ratios.

### v2

@@ 3.4 Long and Loopable Scene Video Generation
Figure 6: Example frames from generated video at the first, 16th, 32th, 48th, 64th and 80th frames generate by a diffusion model that is capable for 16 frames originally. Despite the increasing video length, the visual quality of the panorama remains consistent, demonstrating the effectiveness of our method in generating long videos.
@@
Most existing dynamic scene synthesis approaches like  [25] and  [18] are constrained in limited duration video sequences, as most video diffusion models produce only short clips (typically 16 frames, 2s durantion at 8 FPS). To address this, we extended Offset Shifting Denoising from spatial to temporal dimension and introduced a temporal shifting strategy that enables continuous motion and the generation of videos with much longer durationby regular video diffusion models $\theta$ that have limited frames length capacity $F_{\theta}$. In detail, the latent of long duration video $Z\in\mathbb{R}^{F^{\prime}\times W\times H}$ is divided along temporal dimention $F^{\prime}$ into $n_{F}$ shorter frame clip windows with duration $F_{\theta}$ each, as shown in Figure 5. which can be formulated by:
$Z_{t}=Con_{F}|_{1:n_{F}}(\Phi_{\theta}(t,c,Split_{F}(Z_{t-1},i,t,n_{F})))$ (7)
where $Split_{F}(X,i,j,t,n_{F})$ indicates dividing the panorama video latent $X$ into $n_{F}$ frame clip windows with offset at time $t$ and select the ithi-th one; $Con_{F}(\cdot)$ means collecting latent from divided windows and concat them to form the complete latent. The shifting offset in temporal dimensions also create ’cross step overlap’ between windows analogous to spatial OSD, enabling generation of videos with much longer duration by regular video diffusion models that have limited frame length, while perserving continuous motion and smooth transitions between frames, as shown in Figure 6.
@@
Moreover, by applying similar looping mechanism in our spatiall OSD process to the temporal dimension we can achieve seamless looping video generation. Specifically, instead of the padding windows at the start and the end of the frame sequence, the first and last frame are regarded as connected and windows are allowed to cross the start and end boundary, treating the video frames as a ring, to create a loopable sequence with smooth, continuous motion.
@@
Source Tuning-Free Arbitrary Size 360° Field-of-View Text Only Condition Image Condition Unlimited Video Length Loopable Generation
360DVD [25] CVPR24 $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$
4K4DGen [18] Arxiv24 $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$
Scalecrafter [11] ICLR24 $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$
VividDream [15] Arxiv24 $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$
DynamicScaler - $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@ 4.1 Qualitative Results
Image
Quality
$\uparrow$
Dynamic
Degree
$\uparrow$
Motion
Smoothness
$\uparrow$
Temporal
Flickering
$\uparrow$
Q-Align(V)
$\uparrow$
@@
We evaluate the performance of DynamicScaler against state-of-the-art models. We detail the implementation, experimental settings, comparison methods, and user studies conducted to assess the efficacy of our approach. The proposed DynamicScaler is built on top of existing text-to-video generation model from VideoCrafter2 [4] and image-to-video generation model from VideoCrafter1 [3].
Compared to previous methods, DynamicScaler supports a wider range of settings and can generate videos of infinite length, while other methods are limited to generating only finite-length videos. Our approach generates dynamic panoramas at various resolutions, including both standard and 360° Field-of-View (FoV) panoramas, with support for both text- and image-conditioned inputs, as shown in Table 1. Currently, among the available concurrent works, only 360DVD is capable of generating dynamic panoramas, so we compare our method directly with it. We assess the visual quality by randomly selecting views from 100 generated video cases using random camera positions. Key factors such as image quality, dynamic range, motion smoothness, temporal flickering, and scene richness are evaluated according to the protocols of VBench [12] and CogVideoX [28]. Furthermore, we incorporate an LLM-based visual evaluator, Q-Align [26], to score both image and video quality. As shown in Table 1, DynamicScaler outperforms existing methods, demonstrating superior video quality in handling complex dynamic scenes.
Figure 7: To demonstrates the seamlessness of generated panorama videos, we horizontally concatenat the video frames, showcasing continuity across the left and right boundaries.
Figure 8: Qualitative visualization of ablation study in regular perspective setting. Refer to our project page for results in video format.
@@ 4.2 Quantitative Results
As demonstrated in Figure 1, our approach generates panoramas of arbitrary sizes, supporting rectangular and 360° Field of View (FoV) configurations, offering unparalleled flexibility for diverse applications. As shown in Figure 7, the region near horizontal boundaries of generated frames align seamlessly, ensuring flawless 360° view—an essential feature for immersive environments. This highlights the robustness of our shift denoise technique in both non-360° and 360° settings. Furthermore, our method facilitates the creation of scene-level, long duration, and loopable videos through temporal shifting, without requiring additional training or compromising visual quality.
@@ 4.3 User Studies
We conduct user studies to evaluate videos based on five criteria: graphics quality, frame consistency, left-right continuity, content distribution, and motion patterns. 20 Participants are asked to select the video with the highest quality. Table 2 shows that DynamicScaler performs favorably in all criteria. We use two comparison types: same-case comparison (using the same scene for direct comparisons) and random-case comparison (using different scenes due to limited publicly available samples for comparison). DynamicScaler receives the highest ratings in most criteria as shown in Table  2, especially for continuity, content distribution, and motion patterns, demonstrating its superior ability to generate seamless, dynamic panorama.
@@ 4.4 Ablation Studies
We conducted ablation studies to evaluate the impact of each core component in DynamicScaler under four different configurations: Direct Inference, where the video is generated directly at the target resolution without enhancement techniques, but results in out-of-memory (OOM) due to high VRAM demands; Without OSD (Offset Shifting Denoiser), which excludes the offset spatial denoising mechanism that mitigates multi-scale artifacts; Without GMG (Global Motion Guidance), where global motion guidance is omitted, reducing frame-to-frame motion continuity; and the Full Method, which integrates all components. As shown in Table 3, the Full Method consistently outperforms all other configurations across all evaluation metrics, underscoring the significance of each component. Both OSD and GMG play critical roles in improving image quality, dynamic range, motion smoothness, and temporal coherence, while also minimizing temporal flickering. We also provide quantitative results of ablation study in Figure 8 and our project page.
5 Conclusions
This paper presents DynamicScaler for scalable, coherent panoramic dynamic scene synthesis. By introducing Offset Shifting Denoiser, our approach ensures efficient denoising and consistent boundary transitions. And Global Motion Guidance mechanism maintains local detail and global motion continuity, delivering superior content quality and motion smoothness.
Overall, DynamicScaler outperforms existing methods in scalability and performance, offering a practical, training-free solution for creating high-quality, immersive AR/VR dynamic content across various resolutions and aspect ratios.
Acknowledgments
Supported by the Intelligence Advanced Research Projects Activity (IARPA) via Department of Interior/ Interior Business Center (DOI/IBC) contract number 140D0423C0074. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright annotation thereon. Disclaimer: The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of IARPA, DOI/IBC, or the U.S. Government.